OpenStudy (anonymous):

The population of a local species of beetle can be found using an infinite geometric series where a1 = 880 and the common ratio is . Write the sum in sigma notation, and calculate the sum (if possible) that will be the upper limit of this population.

jimthompson5910 (jim_thompson5910):

The common ratio is missing

OpenStudy (anonymous):

1/4

jimthompson5910 (jim_thompson5910):

r = 1/4 = 0.25

jimthompson5910 (jim_thompson5910):

since |r| < 1, this means that the infinite series does converge to a finite number

jimthompson5910 (jim_thompson5910):

S = a/(1 - r) S = 880/(1 - 0.25) S = 1,173.333333... Round up to get 1,174. That is the upper limit of the population. You would represent this in sigma notation as